OpenStudy (anonymous):

Give a unique example of each: 1. An absolute value equation with two solutions 2. An absolute value equation with one solution 3. An absolute value equation with no solutions

OpenStudy (anonymous):

|2x-7|>10

OpenStudy (anonymous):

sorry, it is |2x-7| = 10

OpenStudy (anonymous):

for 3 it is |x+3| = -2

OpenStudy (anonymous):

for the second one |x| = 0

myininaya (myininaya):

also |x-2|=x+7 works too x-2=x+7 x-2=-x-7 0=9 2x=-5 no solution here x=-5/2 this one works for 2nd one too

myininaya (myininaya):

but |x|=0 is easiest